OpenStudy (anonymous):

solve with row operations x+y=6 x-y=2

OpenStudy (anonymous):

show ur work

OpenStudy (anonymous):

no, x=4 and y=2

OpenStudy (anonymous):

r2 <- r2 - r1 x + y = 6 0 -2y = -4 y = 2 x + 2 =6 x = 4

OpenStudy (anonymous):

thx